A triangular parcel of ground has sides of lengths 725ft, 650ft and 575ft. Find the measure of the largest angle.
Answer by fractalier(6550)

We will apply the law of cosines here.
Realize the largest angle is always opposite the largest side. Let's call it c. Then we write (and solve for angle C)
c^2 = a^2 + b^2 - 2ab cos C
so that
cos C = (c^2 - a^2 - b^2)/-2ab = (725^2 - 650^2 - 575^2)/(-2*650*575) = .3043
and
angle C = 72.3 degrees